# Decision Machines: Enhanced Decision Trees

### v6

@@ 3 Analytical Decision Machines
$\begin{split}h&=\mathrm{\bar{B}}\operatorname{sgn}(Sx-t),\\ p&=\arg\max_{d\in\Delta}\left<d,h\right>,\\ T(x)&=\sum_{i=1}^{L}p[i]v[i],\end{split}$ (7)
where the notation $\Delta$ is the probability simplex defined as $\Delta=\{d\mid d\geq 0,d^{T}\vec{1}=1\}\subset\mathbb{R}^{L}$; the optima $p$ is in $\{0,1\}^{L}$ where its only non-zero element is in the position of maximum in $h\in\mathbb{R}^{L}$; the integer $L$ is the total number of leaves. Initially, the hidden decision $h$ for an input instance $x$ is calculated through the feature selection matrix $S$ and the threshold vector $t$. Subsequently, the most probable exit leaf is determined by maximizing the dot product within the probability simplex $\Delta$. Finally, the ultimate prediction $T(x)$ is derived as a weighted sum of the values associated with each leaf, reflecting their respective probabilities. Clearly, the logical decision machine (5) is of hard attention mechanism[36].
@@
$\begin{split}h&=\mathrm{\bar{B}}\operatorname{sgn}(\mathrm{S}x-t),\\ T(x)&=\sum_{i=1}^{L}\delta(1-h[i])v[i],\end{split}$ (8)
where the function $\delta(x)$ is equal to $1$ if $x=0$ and 00 otherwise. And $\delta(\cdot)$ is activated to identify the exit leaf. Obviously, both $\mathrm{\bar{B}}$ and $\mathrm{S}$ are quite sparse matrices, which may help to understand why sparse or quantized neural networks work. However, the shapes of $\mathrm{\bar{B}}$ and $S$ may restrict the design space of logical decision machines. Additionally, the operators $\operatorname{sgn}(\cdot)$, $\delta(\cdot)$ and $\arg\max(\cdot)$ are non-differentiable, which isolates logical decision machines from deep neural networks.
@@
$F(x)=\sum_{i=1}^{N}w_{i}T_{i}(x)$
where $w_{i}\in\mathbb{R}$ and the tree $T_{i}$ is the instances of (8) for $i=1,2,\cdots,N$. It is simple to evaluate (9) in parallel as the same as the random forest [3]. For example, suppose there are two instances $T_{1}(x)$ and ${T}_{2}(x)$ of 8 as following
$h_{1}=\bar{B}_{1}\operatorname{sgn}(S_{1}x-t_{1}),\,T_{1}(x)=\sum_{i=1}^{L_{1}}\delta(1-h_{1}[i])v_{1}[i],$
$h_{2}=\bar{B}_{2}\operatorname{sgn}(S_{2}x-t_{2}),\,T_{2}(x)=\sum_{i=1}^{L_{2}}\delta(1-h_{2}[i])v_{2}[i],$
we can evaluate them in parallel and obtain the sum of trees $w_{1}T_{1}+w_{2}T_{2}$. In fact, the additive tree $w_{1}T_{1}(x)+w_{2}T_{2}(x)$ is yet another instance of (8) as proven below:
$\begin{split}h&=B\operatorname{sgn}(Sx-t),\\ T(x)&=\sum_{i=1}^{L_{1}+L_{2}}\delta(1-h[i])v[i],\end{split}$ (9)
where $B=\begin{pmatrix}\bar{B}_{1}&0\\ 0&\bar{B}_{2}\end{pmatrix}\in\mathbb{R}^{(L_{1}+L_{2})\times(L_{1}+L_{2}-2)},S=\begin{pmatrix}{S}_{1}\\ {S}_{2}\end{pmatrix}\in\{0,1\}^{(L_{1}+L_{2}-2)\times n},$ $t=\begin{pmatrix}t_{1}\\ t_{2}\end{pmatrix}\in\mathbb{R}^{L_{1}+L_{2}-2},v=\begin{pmatrix}w_{1}v_{1}\\ w_{2}v_{2}\end{pmatrix}\in\mathbb{R}^{L_{1}+L_{2}},x\in\mathbb{R}^{n}$. Here we put the multipliers in the value vectors; the function $\delta(\cdot)$ is activated twice for each input example $x$. By this method, we remove the restriction on the shapes of the matrices $\bar{B}$ and $S$ in (5). We are closer to the key-value attention [29, 34].
In the rest of this section, we will build analytical decision machines to approximate or remove some operators in logical decision machines and take advantages of attention mechanisms to expand the design space of decision machines.
@@ 3.1 Construction
The first refinement is to remove the $\operatorname{sgn}$ (1) in (5) or (8). The exit leaf is identified by the position of the leftmost bit set to $1$ in the product of the false nodes’ bit-vectors , where are assigned to the internal nodes to represents false leaves as $1$ and true leaves as 00 in QuickScorer [25]. We can combine those procedures and remove the $\operatorname{sgn}$ operator in (8) with well-designed matrices $F$ as described below
$\begin{split}h&=F\operatorname{ReLU}(Wx-b),\\ T(x)&=\sum_{i=1}^{L}\delta(1-h[i])v[i],\end{split}$ (10)
where the operator $\operatorname{ReLU}(Wx-b)$ is defined as $\max(0,Wx-b)$; other notation is illustrated in(8). The trouble to remove $\operatorname{sgn}$ in (5) is the probability of zero generated by the affine mapping $Wx-t$.
@@
$T(x)=\sum_{i=1}^{L}\alpha_{i}v_{i},$ (11)
where $(\alpha,\cdots,\alpha_{L})^{T}=\operatorname{softmax}(\frac{K\,(Sx-t)}{\tau})$; the optional $\tau$ is the temperature hyperparameter in $(0,1)$. There are diverse variants of $\operatorname{softmax}$ such as [13, 39, 15, 21]. Here we focus on the positively correlated tests and define our alternative to the $\operatorname{softmax}$ operator in (11) by
@@
where the vector $q=Sx-t$; the real $sim(K_{i},q)_{+}$ is the positive part of the $sim(K_{i},q)$ defined by $\max(0,\left<K_{i},q\right>)$; $S$ is $\sum_{i=1}^{L}sim(K_{i},q)_{+}$; the optional $\alpha$ is a hyperparameter in $(0,1]$. It is a special case of [15].
The third refinement is aimed at realization of the soft decision tree[12]. In [12], each internal node directs instances to its children with probabilities given by a sigmoid gating function. It is equivalent to assign a vector to each branch where left leaves are represented as the sigmoid function $\sigma(\cdot)$; right leaves are represented as $1-\sigma(\cdot)$; and other leaves are represented as $1$. We can derive a matrix $M$ from the template matrix $B$ to replace $\bar{B}$ in (5) as shown below
$M_{ij}=\begin{cases}\sigma(B_{ij}(w_{j}^{T}x+b_{j})),&\text{if $B_{ij}\neq 0$,}\\ 1,&\text{otherwise},\end{cases}$ (12)
where the function $\sigma(x)=\frac{1}{1+\exp(-x)}$; the integer $B_{ij}\in\{1,0,-1\}$ is the $j$th element in the $i$th row of template matrix $B$ in (4); the weight $w_{j}\in\mathbb{R}^{n}$ and the bias $b_{j}\in\mathbb{R}$ are the parameters associated with $j$th branch. The probability $p[i]$ for each leaf $i$ is equal to $\prod_{j=1}^{L-1}M_{ij}$, which is the product of all the non-zero elements in the $i$th row of $M$. With some arithmetic tricks, the output $y$ of the soft decision tree [12] is evaluated by
$\begin{split}p&=\exp(\ln(M)\vec{1}),\\ y&=\sum_{i=1}^{L}p[i]v[i],\end{split}$ (13)
where the matrix $M$ is defined in (12); the vector $\vec{1}$ is a vector of ones with the same row length of $M$; the vector $p$ is the probability distribution of each leaf, i.e.,$\sum_{i=1}^{L}p[i]=1$.
Matrices in (5), (9), (10) and (12) contain diverse hierarchical information.
3.2 Attention-Inspired models
@@
$Attention(Q,K,V)=\operatorname{softmax}(\frac{QK^{T}}{\sqrt{d_{k}}})V,$
where a set of queries are packed together into a matrix $Q$ in order to compute the attention function on these queries simultaneously. We can evaluate the attention output $y_{i}$ for each query $q_{i}$ separately as follows
$y_{i}=\operatorname{softmax}(\frac{q_{i}K^{T}}{\sqrt{d_{k}}})V=\sum_{j=1}^{m}\frac{sim(k_{j},q_{i})}{\sum_{n=1}^{i}sim(k_{n},q_{i})}v_{j},$ (14)
where the term $sim(k_{j},q_{i})=\exp(\frac{\left<k_{j},q_{i}\right>}{\sqrt{d_{k}}})$ represents the similarity between the query $q_{i}$ and the key $k_{j}$; the integer $d_{k}$ is the dimension of the key $q_{i}$. Our algorithm (11) is almost identical to the scaled dot-product attention (14), except for the divided temperature of $\tau$. It is a strong evidence that we can derivate the attention mechanism from the decision tree. We relaxed the binary decision tree in (11) from the perspective of model representation, while the scaled dot-product attention is used to let sequence transduction models dispense with recurrence and convolutions in [34].
As shown before, it is the key matrix $B$ that determines the intrinsic hierarchical structure of (9). We can mask the sub-structure of additive trees by matrix projection. For example, We can mask the second tree in (9):
$\begin{split}h_{B}&=BW^{B}\operatorname{sgn}(Sx-t),\\ T_{B}(x)&=\sum_{i=1}^{L_{1}+L_{2}}\delta(1-h_{B}[i])v[i],\end{split}$
where the matrix $W^{B}=\begin{pmatrix}1_{L_{1}\times(L_{1}-1)}&0\\ 0&0_{L_{2}\times(L_{2}-1)}\end{pmatrix}$; the rest are the same as in (9). Here we projected $B$ of (9) to $\begin{pmatrix}\bar{B}_{1}&0\\ 0&0\end{pmatrix}$. This finding may provide some insights into multi-head attention, which originated in [34] to capture the information from different representation subspaces at different positions

### v7

@@ 3 Analytical Decision Machines
$\begin{split}\mathbf{h}&=\operatorname{sgn}(\mathrm{S}\mathbf{x}-\mathbf{t}),\\ \mathbf{p}&=\arg\max_{\mathbf{d}\in\Delta}\left<\mathbf{d},\mathrm{\bar{B}}\mathbf{h}\right>,\\ T(\mathbf{x})&=\sum_{i=1}^{L}p[i]v[i],\end{split}$ (7)
where the notation $\Delta$ is the probability simplex defined as $\Delta=\{\mathbf{d}\mid\mathbf{d}\geq 0,\mathbf{d}^{T}\vec{1}=1\}\subset\mathbb{R}^{L}$; the optima $\mathbf{p}$ is in $\{0,1\}^{L}$ where its only non-zero element is in the position of maximum in $\mathbf{h}\in\mathbb{R}^{L}$; the integer $L$ is the total number of leaves. Initially, the hidden decision $\mathbf{h}$ for an input instance $\mathbf{x}$ is calculated through the feature selection matrix $\mathrm{S}$ and the threshold vector $\mathbf{t}$. Subsequently, the most probable exit leaf is determined by maximizing the dot product within the probability simplex $\Delta$. Finally, the ultimate prediction $T(\mathbf{x})$ is derived as a weighted sum of the values associated with each leaf, reflecting their respective probabilities. Clearly, the logical decision machine (5) is of hard attention mechanism[39].
@@
$\begin{split}\mathbf{h}&=\mathrm{\bar{B}}\operatorname{sgn}(\mathrm{S}\mathbf{x}-\mathbf{t}),\\ T(\mathbf{x})&=\sum_{i=1}^{L}\delta(1-\mathbf{h}[i])\mathbf{v}[i],\end{split}$ (8)
where the function $\delta(z)$ is equal to $1$ if $z=0$ and 00 otherwise. And $\delta(\cdot)$ is activated to identify the exit leaf. Obviously, both $\mathrm{\bar{B}}$ and $\mathrm{S}$ are quite sparse matrices, which may help to understand why sparse or quantized neural networks work. However, the shapes of $\mathrm{\bar{B}}$ and $\mathrm{S}$ may restrict the design space of logical decision machines. Additionally, the operators $\operatorname{sgn}(\cdot)$, $\delta(\cdot)$ and $\arg\max(\cdot)$ are non-differentiable, which isolates logical decision machines from deep neural networks.
@@
$F(\mathbf{x})=\sum_{i=1}^{N}w_{i}T_{i}(\mathbf{x})$
where $w_{i}\in\mathbb{R}$ and the tree $T_{i}$ is the instances of (8) for $i=1,2,\cdots,N$. It is simple to evaluate (9) in parallel as the same as the random forest [3]. For example, suppose there are two instances $T(\mathbf{x})$ and $\tilde{T}(\mathbf{x})$ of 8 as following
$\bar{\mathbf{h}}=\bar{\mathrm{B}}\operatorname{sgn}(\bar{\mathrm{S}}\mathbf{x}-\bar{\mathbf{t}}),\,{T}_{1}(\mathbf{x})=\sum_{i=1}^{L_{1}}\delta(1-\bar{\mathbf{h}}[i])\bar{\mathbf{v}}[i],$
$\tilde{\mathbf{h}}=\tilde{\mathrm{B}}\operatorname{sgn}(\tilde{\mathrm{S}}\mathbf{x}-\tilde{\mathbf{t}}),\,{T}_{2}(\mathbf{x})=\sum_{i=1}^{L_{2}}\delta(1-\tilde{\mathbf{h}}[i])\tilde{\mathbf{v}}[i],$
we can evaluate them in parallel and obtain the sum of trees $w_{1}\bar{T}+w_{2}\tilde{T}$. In fact, the sum of the decision trees is yet another instance of (8) as proven below:
$\begin{split}\mathbf{h}&=\mathrm{B}\operatorname{sgn}(\mathrm{S}\mathbf{x}-\mathbf{t}),\\ T(\mathbf{x})&=\sum_{i=1}^{L_{1}+L_{2}}\delta(1-\mathbf{h}[i])\mathbf{v}[i],\end{split}$ (9)
where $\mathrm{S}=\begin{pmatrix}\bar{\mathrm{S}}\\ \tilde{\mathrm{S}}\end{pmatrix},\mathbf{t}=\begin{pmatrix}\bar{\mathbf{t}}\\ \tilde{\mathbf{t}}\end{pmatrix},\mathrm{B}=\begin{pmatrix}\bar{\mathrm{B}}&0\\ 0&\tilde{\mathrm{B}}\end{pmatrix},\mathbf{v}=\begin{pmatrix}w_{1}\bar{\mathbf{v}}\\ w_{2}\tilde{\mathbf{v}}\end{pmatrix}$.Here we put the multipliers in the value vectors; the function $\delta(\cdot)$ is activated twice for each input example $\mathbf{x}$. By this method, we remove the restriction on the shapes of the matrices $\bar{\mathrm{B}}$ and $\mathrm{S}$ in (5). We are closer to the key-value attention [29, 37].
In the rest of this section, we will construct analytical decision machines to approximate or remove some operators in logical decision machines and take advantages of attention mechanisms to expand the design space of decision machines.
@@ 3.1 Construction
The first refinement is to remove the $\operatorname{sgn}$ function defined in 1. The exit leaf is identified by the position of the leftmost bit set to $1$ in the product of the false nodes’ bit-vectors, which are assigned to the internal nodes to represents false leaves as $1$ and true leaves as 00 in QuickScorer [25]. Similarly, it is able to identify the exit leaf by the position of the rightmost bit set to $1$ in the product of the true nodes’ bit-vectors if the false nodes and true nodes are represented as 00 and $1$ in the bitvectors respectively. The $\operatorname{sgn}(\cdot)$ operator in (7) is removed by combining both procedures with well-designed matrices $\mathrm{F}$ as described below
$\begin{split}\mathbf{h}&=\mathrm{F}\operatorname{ReLU}(\mathrm{W}\mathbf{x}-\mathbf{b}),\\ T(\mathbf{x})&=\mathbf{v}[\arg\max(\mathbf{h})],\end{split}$ (10)
where the operator $\operatorname{ReLU}(\mathrm{W}\mathbf{x}-\mathbf{b})$ is defined as $\max(0,\mathrm{W}\mathbf{x}-b)$; $\arg\max(\cdot)$ returns the smallest index of the maximum value in the input vector $\mathbf{h}$.
@@
$T(\mathbf{x})=\sum_{i=1}^{L}\alpha_{i}\mathbf{v}_{i},$ (11)
where $(\alpha,\cdots,\alpha_{L})^{T}=\operatorname{softmax}(\frac{\mathrm{K}\,(\mathrm{S}\mathbf{x}-\mathbf{t})}{\tau})$; the optional $\tau$ is the temperature hyperparameter in $(0,1)$. There are diverse variants of $\operatorname{softmax}$ such as [13, 42, 15, 21]. Here we focus on the positively correlated tests and define our alternative to the $\operatorname{softmax}$ operator in (11) by
@@
where the vector $q=Sx-t$; the real $sim(K_{i},q)_{+}$ is the positive part of the $sim(K_{i},q)$ defined by $\max(0,\left<K_{i},q\right>)$; $\mathrm{S}$ is $\sum_{i=1}^{L}sim(K_{i},q)_{+}$; the optional $\alpha$ is a hyperparameter in $(0,1]$. It is a special case of [15].
The third refinement is aimed at realization of the soft decision tree[12]. In [12], each internal node directs instances to its children with probabilities given by a sigmoid gating function. It is equivalent to assign a vector to each branch where left leaves are represented as the sigmoid function $\sigma(\cdot)$; right leaves are represented as $1-\sigma(\cdot)$; and other leaves are represented as $1$. We can derive a matrix $\mathrm{M}$ from the template matrix $\mathrm{B}$ to replace $\bar{\mathrm{B}}$ in (5) as shown below
$\mathrm{M}_{ij}=\begin{cases}\sigma(\mathrm{B}_{ij}(\mathbf{w}_{j}^{T}\mathbf{x}+\mathbf{b}_{j})),&\text{if $\mathrm{B}_{ij}\neq 0$,}\\ 1,&\text{otherwise},\end{cases}$ (12)
where the function $\sigma(z)=\frac{1}{1+\exp(-z)}$; the integer $\mathrm{B}_{ij}\in\{1,0,-1\}$ is the $j$th element in the $i$th row of template matrix $\mathrm{B}$ in (4); the weight $\mathbf{w}_{j}\in\mathbb{R}^{n}$ and the bias $b_{j}\in\mathbb{R}$ are the parameters associated with $j$th branch. The probability $p[i]$ for each leaf $i$ is equal to $\prod_{j=1}^{L-1}\mathrm{M}_{ij}$, which is the product of all the non-zero elements in the $i$th row of $\mathrm{M}$. With some arithmetic tricks, the output $y$ of the soft decision tree [12] is evaluated by
$\begin{split}\mathbf{p}&=\exp(\ln(\mathrm{M})\vec{1}),\\ y&=\sum_{i=1}^{L}\mathbf{p}[i]\mathbf{v}[i],\end{split}$ (13)
where the matrix $M$ is defined in (12); the vector $\vec{1}$ is a vector of ones with the same row length of $\mathrm{M}$; the vector $\mathbf{p}$ is the probability distribution of each leaf, i.e.,$\sum_{i=1}^{L}p[i]=1$.
Matrices in (5), (9), (10) and (12) contain diverse hierarchical information, which makes it distinct from the nested logit model and hierarchical softmax [32, 30, 31].
3.2 Attention Models
@@
$Attention(\mathrm{Q},\mathrm{K},\mathrm{V})=\operatorname{softmax}(\frac{\mathrm{Q}\mathrm{K}^{T}}{\sqrt{d_{k}}})\mathrm{V},$
where a set of queries are packed together into a matrix $\mathrm{Q}$ in order to compute the attention function on these queries simultaneously. We can evaluate the attention output $\mathbf{y}_{i}$ for each query $\mathbf{q}_{i}$ separately as follows
$\mathbf{y}_{i}=\operatorname{softmax}(\frac{\mathbf{q}_{i}\mathrm{K}^{T}}{\sqrt{d_{k}}})\mathrm{V}=\sum_{j=1}^{m}\frac{sim(\mathbf{k}_{j},\mathbf{q}_{i})}{\sum_{n=1}^{i}sim(\mathbf{k}_{n},\mathbf{q}_{i})}\mathbf{v}_{j},$ (14)
where the term $sim(\mathbf{k}_{j},\mathbf{q}_{i})$ is defined as $\exp(\left<\mathbf{k}_{j},\mathbf{q}_{i}\right>)$ to measure the similarity between the query $\mathbf{q_{i}}$ and the key $\mathbf{k_{j}}$; the integer $d_{k}$ is the dimension of the key $\mathbf{q_{i}}$. Our algorithm (11) is almost identical to the scaled dot-product attention (14), except for the divided temperature of $\tau$. It is a strong evidence that we can derivate the attention mechanism from the decision tree. We relaxed the binary decision tree in (11) from the perspective of model representation, while the scaled dot-product attention is used to let sequence transduction models dispense with recurrence and convolutions in [37].
As shown before, it is the key matrix $\mathrm{B}$ that determines the intrinsic hierarchical structure of (9). We can mask the sub-structure of additive trees by matrix projection. For example, We can mask the second tree in (9):
$\begin{split}h_{B}&=\mathrm{B}W^{B}\operatorname{sgn}(\mathrm{S}\mathbf{x}-\mathbf{t}),\\ T_{B}(x)&=\sum_{i=1}^{L_{1}+L_{2}}\delta(1-h_{B}[i])v[i],\end{split}$
where the matrix $W^{B}=\begin{pmatrix}1_{L_{1}\times(L_{1}-1)}&0\\ 0&0_{L_{2}\times(L_{2}-1)}\end{pmatrix}$; the rest are the same as in (9). Here we projected $\mathrm{B}$ of (9) to $\begin{pmatrix}\bar{B}_{1}&0\\ 0&0\end{pmatrix}$. This finding may provide some insights into multi-head attention, which originated in [37] to capture the information from different representation subspaces at different positions